# SwitchHead: Accelerating Transformers with Mixture-of-Experts Attention

### v2

@@ 3.1 Which Projections Require an MoE?
Table 1: Performance of SwitchHead with $E=5$ experts and $H=2$ heads. Different projections are either experts or fixed for the given head. Parameter-matched baseline with $H=10$ and $H=2$ are shown. Models sorted by perplexity. 47M parameters models on Wikitext 103.
Model $n_{\text{heads}}$ V expert K expert Q expert O expert Perplexity
@@
3.2 Comparing with MoA
The method most related to ours is the so-called Mixture of Attention Heads, or MoA (Zhang et al., 2022). They use a selection mechanism to choose active attention heads from a set of experts. However, they have a single set of $K$ and $V$ projections shared between experts; thus, acceleration is possible. However, in the original paper, the authors use a high number of selected heads (8-16) which seems necessary to achieve good performance. Thus, the resource reductions are moderate. Moreover, MoA uses three different regularizers, which have to be tuned independently.
We compare our method with our reimplementation of MoA with a different number of selected heads. Given the complexity of tuning its regularization coefficients, we take them directly from Zhang et al. (2022). For a fair comparison, we also integrated the non-competitive selection mechanism of Csordás et al. (2023) into MoA. The results are shown in Table 2. Similarly to our method, we found that with non-competitive selection, no regularization is required, and the predictive performance usually is superior to the original formulation. However, it still underperforms our method given a similar computation and memory budget.
Table 2: Performance of SwitchHead compared to different MoA variants. "sel. activation" denotes the activation function used for selecting the experts. the MoA can outperform the baseline, but only at a price of using significantly more computing and memory. Also, SwitchHead outperforms the baseline dense Transformer. Results are on Wikitext 103.
Model sel. activation $n_{\text{heads}}$params Perplexity MACs Mem (floats)
MoA sigmoid 8 47M 12.13 390.2M 2.6M
MoA sigmoid 6 47M 12.16 306.8M 1.9M
SwitchHead sigmoid 2 47M 12.27 170.4M 0.8M
Transformer XL - 10 47M 12.31 453.4M 3.5M
MoA sigmoid 4 47M 12.39 223.5M 1.3M
MoA softmax 4 47M 12.60 223.5M 1.3M
MoA softmax 6 47M 12.64 306.8M 1.9M
MoA sigmoid 2 47M 12.65 140.1M 0.7M
MoA softmax 8 47M 12.77 390.2M 2.6M
MoA softmax 2 47M 12.84 140.1M 0.7M
MoA softmax 8 262M 9.50 2.9G 9.9M
SwitchHead sigmoid 2 262M 9.55 2.0G 2.9M
MoA sigmoid 8 262M 9.56 2.9G 9.9M
MoA sigmoid 12 262M 9.58 4.1G 14.7M
Transformer XL - 16 262M 9.66 5.4G 21.0M
MoA softmax 12 262M 9.68 4.1G 14.7M
MoA softmax 4 262M 9.69 1.7G 5.1M
MoA sigmoid 4 262M 9.77 1.7G 5.1M
MoA softmax 2 262M 9.87 1.1G 2.7M
MoA sigmoid 2 262M 10.02 1.1G 2.7M
3.3 Performance on Different Datasets
We test our methods on a diverse set of language modeling datasets, including C4 (Raffel et al., 2020), Enwik8 (Hutter, 2006), peS2o (Soldaini & Lo, 2023), at two different scales: a 47M and a 262M parameters. The results are shown in Tab. 3. We compare our models to two baselines: one with the same number of heads as the total number of experts ($H\cdot E$) of the SwitchHead models, and the other has the same number of heads as the number of active attention matrices ($H$) as our models. Our models always closely match the performance of the full, many-head baseline with the fraction of memory and compute requirements. Importantly, our method also achieves a wall-clock speedup, enough to accelerate the entire training pipeline by a factor of around 1.5 (see Appendix A.4 for more details). This confirms the competitiveness of our method.
Table 3: Performance of SwitchHead compared to baselines on different datasets with different model sizes. It can be seen that the predictive performance of our SwitchHead model is comparable to the baselines, and is always better than the baseline with an equal number of heads. Perplexity is shown for Wikitext 103, C4 and peS2o datasets, and bits/character (bpc) for Enwik8.
@@ 3.3 Performance on Different Datasets
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchHead C4 2 47M 22.49 202.5M 0.8M
Transformer XL C4 10 47M 22.63 453.4M 3.5M
Transformer XL C4 2 47M 23.71 453.4M 1.4M
SwitchHead C4 4 262M 16.38 2.4G 5.6M
Transformer XL C4 16 262M 16.58 5.4G 21.0M
Transformer XL C4 4 262M 17.09 5.4G 8.4M
SwitchHead Wikitext 103 2 47M 12.31 170.4M 0.8M
Transformer XL Wikitext 103 10 47M 12.32 453.4M 3.5M
Transformer XL Wikitext 103 2 47M 12.73 453.4M 1.4M
SwitchHead Wikitext 103 2 262M 9.77 2.0G 2.9M
Transformer XL Wikitext 103 16 262M 9.80 5.4G 21.0M
Transformer XL Wikitext 103 2 262M 10.09 5.4G 6.3M
Transformer XL peS2o 10 47M 12.83 453.4M 3.5M
SwitchHead peS2o 2 47M 12.84 202.5M 0.8M
Transformer XL peS2o 2 47M 13.37 453.4M 1.4M
Transformer XL peS2o 16 262M 9.78 5.4G 21.0M
SwitchHead peS2o 4 262M 9.86 2.4G 5.6M
Transformer XL peS2o 4 262M 10.11 5.4G 8.4M
SwitchHead Enwik8 2 41M 1.10 709.3M 2.8M
Transformer XL Enwik8 8 41M 1.10 1.6G 10.5M
Transformer XL Enwik8 2 41M 1.13 1.6G 4.2M
3.3.1 SwitchAll
The goal of achieving more resource-efficient Transformers includes reducing the resource requirements of both the MLP and the attention layers. Csordás et al. (2023) proposed a parameter-efficient MoE method to accelerate the MLP layers. However, it remains unclear whether it can be efficiently combined with our SwitchHead, or can have some negative interaction effect if combined in a "SwitchAll", where every layer is MoE-based.
In order to verify this, we take the architecture proposed by Csordás et al. (2023) without any hyperparameter change and replace the attention layer with SwitchHead. The hyperparameters for the attention are directly taken from the experiments shown in Tab. 3. The results are shown in Tab. 4. The combined, fully-MoE model often outperforms the dense baselines for each dataset and model size considered, except in the case of the 259M parameter model on the C4 dataset.
Table 4: Performance of SwitchAll (SwitchHead + $\sigma$-MoE (Csordás et al., 2023)) on different datasets and model sizes. Our SwitchAll model is close or better compared to the baselines.
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchAll Wikitext 103 2 47M 12.17 170.4M 0.8M
Transformer XL Wikitext 103 10 47M 12.32 453.4M 3.5M
Transformer XL Wikitext 103 16 262M 9.80 5.4G 21.0M
SwitchAll Wikitext 103 4 259M 9.81 2.4G 5.6M
SwitchAll C4 2 47M 22.09 202.5M 0.8M
Transformer XL C4 10 47M 22.63 453.4M 3.5M
SwitchAll C4 4 259M 16.45 2.4G 5.6M
Transformer XL C4 16 262M 16.58 5.4G 21.0M
SwitchAll peS2o 2 47M 12.56 202.5M 0.8M
Transformer XL peS2o 10 47M 12.83 453.4M 3.5M
Transformer XL peS2o 16 262M 9.78 5.4G 21.0M
SwitchAll peS2o 4 259M 9.86 2.4G 5.6M
4 RoPE Positional Encodings
All of our experiments so far have used a Transformer XL model. Thus, it remains unclear whether SwitchHead is specific to this model or can be also used with other attention methods. As an alternative, we consider RoPE positional encodings Su et al. (2021) without the XL cache (thus, the attention matrices are square). We test these models on Wikitext 103. The results are shown in Tab. 5. Our method also performs well in this case.
@@ 4 RoPE Positional Encodings
Table 5: Performance of SwitchHead compared to dense baseline on Wikitext 103, using RoPE positional encoding instead of Transformer XL.
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchHead (RoPE) Wikitext 103 2 45M 12.75 285.6M 1.3M
Transformer (RoPE) Wikitext 103 10 45M 12.78 560.9M 6.1M
Transformer (RoPE) Wikitext 103 2 45M 12.96 560.9M 1.9M
SwitchHead (RoPE) Wikitext 103 4 243M 10.00 4.2G 18.4M
Transformer (RoPE) Wikitext 103 16 244M 10.17 6.4G 37.7M
Transformer (RoPE) Wikitext 103 2 244M 10.26 6.4G 8.4M
5 Analysis
In order to see how the network uses the attention heads, we trained a small, 6-layer, 8-head Transformer on ListOps (Nangia & Bowman, 2018; Csordás et al., 2022). The reason for this choice is that small, algorithmic tasks tend to be more interpretable compared to language models. We also train a parameter-matched, 2-head SwitchHead model. Both models achieve around 95% accuracy on a held-out IID validation set, in contrast to the dense 2-head model, which saturates around 80%. Note that ListOps is a classification task and does not use autoregressive masking.
Following Csordás et al. (2022), we visualize the maximum of attention heads for each layer, both for the standard Transformer (Fig. 2a) and SwitchHead (Fig. 2b). The attention maps are qualitatively similar. Note that the initialization and the learning dynamics are different for the two models, thus the overlap would not be perfect even with the same type of model. We show all the attention maps for both models in Fig. 4 and 3 in the Appendix.
In addition, we visualize individual attention heads for the SwitchHead model. An example is shown in Fig. 2c. In addition to the attention map, we show the weight of the selected experts for both the value and output projection (denoted by V and O, respectively, on the sides of the attention map). Often it is possible to interpret the selection weights: here, the output experts specialize according to different operations, while the input ones distinguish numbers and closed parentheses. The attention map itself appears to distribute information about contiguous chunks of numbers. Similar plots for all heads are shown in Fig. 5 in the Appendix.
(a) Transformer, Layer 3
(b) SwitchHead Layer 3
(c) SwitchHead, Layer 3, Head 1
Figure 2: An attention map of the (a) standard Transformer and (b) SwitchHead. The maximum of all heads in the given layer are shown. (c) A head of SwitchHead. On the left side of the attention plot, the selection weights of the output projection expert are shown. Similarly, at the bottom, the selection weights of the value experts are visible. In the selection maps, dark blue always corresponds to 1, while white is 0. The scale shown on the right is only for the attention.
The attention maps of the language models are difficult to interpret. However, we visualized the attention maps of the 47M parameter Transformer XL and the SwitchHead model from Tab. 3. We found them to be qualitatively similar. We also identified induction heads (Olsson et al., 2022) in both models, some examples shown for SwitchHead in Fig. 6a and for Transformer in Fig. 6b in the appendix. Other typical vertical line-lined attention patterns are shown in Fig. 6c and 6d.
6 Related Work
The method most closely related to ours is MoA (Zhang et al., 2022), which introduces a MoE style attention. It defines each attention head as an expert but shares the key and value projections between them. Unlike in our case, each of the selected experts requires a separate attention matrix, which significantly increases its memory usage. Due to the use of a competitive softmax-based activation function in the selection network, it requires complex regularization to prevent expert collapse. In the original formulation, the number of active heads is high. We also confirmed in our experiments that MoA needs many attention heads to match the performance of the dense baseline (see Sec. 3.2), and it is only possible to do so with a significantly higher resource budget than our method.
Nguyen et al. (2022) analyze the attention matrices, and they conclude that they are usually low rank. Motivated by this, the authors construct a few (e.g., 2) "global attention matrices", and they compute each local matrix for specific heads by a weighted average of those. However, they average the logits, not the final matrix, so each individual head-specific matrix has to be computed. This means that in the best case, they can only save half of the computation associated with the attention matrix because the readout (Eq. 3) is still needed. For the same reason, memory savings are also low. The authors also use sampling of the attention matrices.
Peng et al. (2020) proposes to reweight the contribution of each head by a gating function. However, they only reduce the number of total attention heads by one, presumably to compensate for the parameters used by the selection logic. Their goal was not to reduce resource usage but to have better predictive performance, which they achieve. They use a softmax-based competitive selection mechanism. To avoid collapse, the gating function is trained only in some steps.
Csordás et al. (2023) introduce the non-competitive $\sigma$-MoE method that we also use for our attention mechanism. However, the authors focus on accelerating the MLPs and not the attention. More broadly, Shazeer et al. (2017) introduces sparsely-gated mixture of experts in LSTM (Hochreiter & Schmidhuber, 1997) networks. Fedus et al. (2021) introduces Mixture of Experts in Transformers. Lepikhin et al. (2021) trains a MoE-based LLM, and Clark et al. (2022) analyzes the scaling laws of MoE models. Lewis et al. (2021) introduces an alternative method for preventing collapse.
Dao et al. (2022) provides a hardware-aware CUDA implementation of the entire attention layer, which avoids storing the attention matrix. By saving memory bandwidth in this way, they achieve a significant wall clock time speedup, despite that the attention matrix should be recomputed in the backward pass. This is orthogonal to our method and they can be combined for further acceleration.
7 Conclusion
On a wide range of language modeling datasets with different model sizes, our novel Mixture-of-Experts-based attention method called SwitchHead achieves performance on par with parameter-matched dense counterparts, but with only a fraction of the computational cost and memory usage. SwitchHead drastically reduces the number of attention matrices that have to be computed, by using MoE for the value and output projections. Our method is stable and does not need additional regularization to prevent degenerate solutions (a well-known practical issue in many existing MoE models). Our method can also be successfully combined with MoE MLP layers, to obtain "SwitchAll" where every layer of the Transformer is MoE-based, achieving a huge reduction in resource requirements.
Acknowledgements
Appendix A Appendix
A.1 Resource Requirements of MoA
The resource requirements of MoA (Peng et al., 2020) are very similar to those of Transformer XL (see Sec. 2.3 for more details), except that it uses a single shared ${\bm{K}}$ and ${\bm{V}}$ for each head.
$N_{\text{MAC}}^{\text{MoA}}$ $=(2H+2)Td_{\text{head}}d_{\text{model}}+2HCT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}$ (14)
$N_{\text{mem}}^{\text{MoA}}$ $=(2H+2)Td_{\text{head}}+2HCT^{2}+2CTd_{\text{head}}$ (15)
A.2 Hyperparameters
@@
We train all our models with Adam optimizer (Kingma & Ba, 2015), with a batch size of 64, a learning rate of 0.00025, and gradient clipping with a maximum norm of $\kappa$. Large models ($>200K$ parameters) use a learning rate warm-up of 4k steps. All models, except the SwitchAll model, use a dropout on the MLP layers, $0.1$ for the small models and $0.2$ for the large ones. Detailed hyperparameters are shown in the Tab. 6. $\sigma$-MoE related hyperparameters for the SwitchAll models are identical to those of Csordás et al. (2023). For Transformer XL models, we always use a single additional chunk of context, both in training and validation time. $d_{\text{head}}$ and $d_{\text{ff}}$ are derived in a systematic way, see Sec. 3 for more details.
@@
Transformer XL Wikitext 103 10 47M 41 2053 - - 256 16 0.1
Transformer XL Wikitext 103 2 47M 205 2053 - - 256 16 0.1
@@
Transformer XL Wikitext 103 16 262M 64 4110 - - 512 18 0.25
Transformer XL Wikitext 103 2 262M 512 4110 - - 512 18 0.25
@@
Transformer XL Enwik8 8 41M 64 2053 - - 512 12 0.25
Transformer XL Enwik8 2 41M 256 2053 - - 512 12 0.25
@@
A.2.1 A Note on Choosing $n_{\text{heads}}$
Our preliminary experiments showed that a single head is usually not enough to match the performance of the baseline network, but two heads usually work well. Because of this, we always start by training a model with $n_{\text{heads}}=2$ and increase it to $n_{\text{heads}}=4$ if it does not match the performance of the baseline. We have not experimented with any other $n_{\text{heads}}$.
A.3 A Note on the Parameter Count of the SwitchAll
It can be seen in Tab. 4 that the parameter count of the SwitchAll models is often less than that of their dense counterparts. The reason is that we normally compensate for the final difference in the number of parameters by increasing $d_{\text{ff}}$ (see Sec. 3 for details of the parameter matching). However, that can only be done in a very coarse-grained way with $\sigma$-MoE: the size of all experts must be increased at once, and the CUDA kernel supports only sizes of multiple of 4. Therefore, increasing the size of the experts would add too many parameters and the model would outgrow the baseline. For this reason, we simply keep the hyperparameters for Csordás et al. (2023) and combine them with our SwitchHead configuration from Tab. 3.
A.4 Wall-Clock Time Estimation
In all of our tables, we report the number of multiply-accumulate (MAC) operations following Zhang et al. (2022). The reason for this is that the actual wall-clock time is highly implementation and hardware-dependent. Nevertheless, we measured the runtime and total memory usage of our entire training pipeline (including the feedforward layer) to demonstrate that our current (suboptimal) implementation is already capable of providing wall-clock-time acceleration. We show the results in Tab. 7. The measurements are taken on identical hardware with the same implementation (including for the attention core), the only difference being the MoE-based projections for the attention. It can be seen that for both scales, our method trains around 1.5 times faster, while using 61%-67% as much memory as the baseline. Note that these measurements also include the MLP layers, the optimizer, and the gradient synchronization in the case of multi-GPU training.
Model Size ms/iteration Rel. iter. time RAM/GPU Rel. Mem. #GPUs GPU type
Trafo. XL 47M 473ms/iter 1.0 20.5G 1.0 1 RTX 3090
SwitchHead 342ms/iter 0.72 13.5G 0.65
Trafo. XL 262M 670ms/iter 1.0 20.5G 1.0 8 V100
SwitchHead 442ms/iter 0.65 12.5G 0.61
A.5 Visalizing all Attention Heads
@@
n8HAOVy+T+XLJfLS4ZhTiRJgqIoVBRFIoric47jPnmeB1mW/9rr9ZpSSn3Lsmir1fJZlqWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ A.3 The Importance of Different Projections
Table 6: Performance of SwitchHead with $E=5$ experts and $n_{\text{heads}}=2$ heads. Different projections are either experts or fixed for the given head. Columns V, K, Q, and O show whether the given projection is an expert. Parameter-matched baseline with $n_{\text{heads}}=10$ and $n_{\text{heads}}=2$ are shown. Models sorted by perplexity. 47M parameters models on Wikitext 103.
Model $n_{\text{heads}}$ V K Q O Perplexity $\downarrow$
@@
A.4 RoPE Positional Encodings
All of our experiments in the main paper have used a Transformer XL model. Thus, it remains unclear whether SwitchHead is specific to this model or can be also used with other attention methods. As an alternative, we consider RoPE positional encodings [28] without the XL cache (thus, the attention matrices are square). This is the standard setup used by modern language models, such as all versions of Llama [27]. We tested these models in Wikitext 103 and C4. The results are shown in Tab. 7, and zero-shot performance on downstream tasks in Tab. 8. This shows that SwitchHead also performs well in the standard setup and is not tied to Transformer XL.
Table 7: Perplexity of SwitchHead compared to dense baseline, using RoPE positional encoding and no XL cache. Memory usage is specified in number of floats. Models sorted by perplexity.
@@ A.4 RoPE Positional Encodings
Dataset #total params Model $n_{\text{heads}}$ ppl $\downarrow$ MACs Memory
Wikitext 103 45M SwitchHead 2 12.75 285.6M 1.3M
Transformer 10 12.78 560.9M 6.1M
Transformer 2 12.96 560.9M 1.9M
244M SwitchHead 4 10.00 4.2G 18.4M
Transformer 16 10.17 6.4G 37.7M
Transformer 2 10.26 6.4G 8.4M
C4 45M SwitchHead 2 23.69 285.6M 1.3M
Transformer 10 23.79 560.9M 6.1M
244M SwitchHead 4 16.41 4.2G 18.4M
Transformer 16 16.35 6.4G 37.7M
@@
Table 8: Zero-shot task performance of SwitchHead using RoPE positional encodings and no XL cache, trained on C4 dataset, compared to dense Transformer baseline with matched number of parameters.
Model #total params ppl $\downarrow$ Lambada $\uparrow$ BLiMP $\uparrow$ CBT $\uparrow$
SwitchHead 45M $23.69$ $20.9\%$ $77.3\%$
Transformer 45M $23.76$ $20.3\%$ $73.8\%$
SwitchHead MAC-matched 54M $22.18$ $22.6\%$ $77.4\%$
SwitchHead Shared selection 45M $23.63$ $20.3\%$ $76.0\%$
SwitchHead 243M $16.41$ $30.5\%$ $79.9\%$ $83.8\%$
Transformer 243M $16.35$ $29.8\%$ $76.1\%$ $83.9\%$
SwitchHead MAC-matched 314M $15.63$ $30.5\%$ $80.5\%$ $84.6\%$
SwitchHead Shared selection 243M $16.59$ $28.1\%$ $79.1\%$ $83.7\%$
A.5 Hyperparameters
@@
We train all our models with Adam optimizer [40], with a batch size of 64, a learning rate of 0.00025, and gradient clipping with a maximum norm of $\kappa$. Large models ($>200K$ parameters) use a learning rate warm-up of 4k steps. All models, except the SwitchAll model, use a dropout on the MLP layers, $0.1$ for the small models and $0.2$ for the large ones. Detailed hyperparameters are shown in the Tab. 9. $\sigma$-MoE related hyperparameters for the SwitchAll models are identical to those of Csordás et al. [17]. For Transformer XL models, we always use a single additional chunk of context, both in training and validation time. $d_{\text{head}}$ and $d_{\text{ff}}$ are derived in a systematic way, see Sec. 3 for more details.
@@
Transformer 10 47M 41 2053 - - 256 16 0.1
Transformer 2 47M 205 2053 - - 256 16 0.1
@@
Transformer 16 262M 64 4110 - - 512 18 0.25
Transformer 2 262M 512 4110 - - 512 18 0.25
@@
Transformer 8 41M 64 2053 - - 512 12 0.25
Transformer 2 41M 256 2053 - - 512 12 0.25
@@
A.6 A Note on the Parameter Count of the SwitchAll
It can be seen in Tab. 3 that the parameter count of the SwitchAll models is often less than that of their dense counterparts. The reason is that we normally compensate for the final difference in the number of parameters by increasing $d_{\text{ff}}$ (see Sec. 3 for details of the parameter matching). However, that can only be done in a very coarse-grained way with $\sigma$-MoE: the size of all experts must be increased at once, and the CUDA kernel supports only sizes of multiple of 4. Therefore, increasing the size of the experts would add too many parameters and the model would outgrow the baseline. For this reason, we simply keep the hyperparameters for Csordás et al. [17] and combine them with our SwitchHead configuration from Tab. 2.
A.7 Visalizing all Attention Heads
@@
A.8 Compute Requirements
We report the compute used for our experiments, including the GPU type, count (the number of GPUs used per experiment, and not the total in the machine), and the runtime in “hh:mm” format in Tab. 10. We report the total number of CPUs ($N_{\text{CPU}}$) and RAM because they are shared between concurrent runs. Note that most of the experiments were done prior to the much faster, Triton-based kernel implementation. Because of this, the runtimes appear longer for SwitcHead compared to the baseline. For timing benchmarks with our new kernel, see Tab. 5.
Note that we only report the resources used for the paper here. We estimate that the total cost of the failed experiments and preliminary runs is around 10 times higher than this.
Table 10: Training hardware information for the experiments reported in the paper
Model #params Dataset $G$ GPU Type $N_{\text{GPU}}$ $N_{\text{CPU}}$ RAM Duration
SwitchAll 259M C4 4 V100-32GB-LS 8 40 503G 24:06
SwitchAll 259M peS2o 4 V100-32GB-LS 8 40 503G 30:00
SwitchAll 259M Wikitext 103 4 RTX 4090 4 24 251G 22:58
SwitchAll 47M C4 2 RTX 3090 1 24 220G 22:14
SwitchAll 47M peS2o 2 RTX 3090 1 24 220G 22:49
SwitchAll 47M Wikitext 103 2 RTX 3090 1 24 251G 6:03
SwitchHead 243M Wikitext 103 4 V100-32GB 4 40 503G 147:09
SwitchHead 262M C4 4 V100-32GB-LS 8 40 503G 26:38
SwitchHead 262M peS2o 4 V100-32GB-LS 8 40 503G 27:43
SwitchHead 262M Wikitext 103 2 V100-32GB 4 40 503G 31:42
SwitchHead 41M Enwik8 2 V100-32GB 1 40 503G 13:45
SwitchHead 45M Wikitext 103 2 RTX 3090 1 24 251G 17:28
SwitchHead 47M C4 2 V100-32GB 1 40 503G 15:36
SwitchHead 47M peS2o 2 V100-32GB 1 40 503G 16:17
SwitchHead 47M Wikitext 103 2 RTX 3090 1 24 251G 13:09
Transformer 262M C4 4 V100-32GB 8 40 503G 11:55
Transformer 262M C4 16 V100-32GB-LS 8 40 503G 20:21
Transformer 262M peS2o 4 V100-32GB 8 40 503G 17:08
Transformer 262M peS2o 16 V100-32GB-LS 8 40 503G 25:56
Transformer 262M Wikitext 103 2 P100-16GB 8 12 62G 0:00
Transformer 262M Wikitext 103 16 A100-80GB 2 64 503G 31:51
Transformer 41M Enwik8 2 RTX 3090 1 24 220G 15:38
Transformer 41M Enwik8 8 V100-32GB-LS 2 40 503G 16:04
Transformer 47M C4 2 V100-32GB 1 40 503G 10:29
Transformer 47M C4 10 V100-32GB 1 40 503G 16:57
Transformer 47M peS2o 2 V100-32GB 1 40 503G 11:07
Transformer 47M peS2o 10 V100-32GB 1 40 503G 17:55
Transformer 47M Wikitext 103 2 V100-32GB 1 40 503G 10:06
Transformer 47M Wikitext 103 10 V100-32GB 1 40 503G 18:51
Transformer (RoPE) 244M Wikitext 103 16 RTX 3090 4 24 251G 30:30
Transformer (RoPE) 45M Wikitext 103 10 V100-32GB 1 40 503G 15:30